OpenStudy (anonymous):

What is the volume of a pyramid with base area B = 9 ft2 and height h = 11 ft? A. 33 ft3 B. 49.5 ft3 C. 66 ft3 D. 99 ft3 @paki

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

V = (L * W * H) / 3 H = 11 base area = L * W....so LW = 9 sub V = (9 * 11) / 3 V = 99/3 V = 33 ft^3 I think I did that right...you might want to get a second opinion